# Overcoming data challenges through enriched validation and targeted sampling to measure whole-person health in electronic health records

### v4

@@ 2.4 Design: Targeted Sampling to Select Informative Patients for Validation
Based on our analysis plan, we considered various sampling designs to select the most informative $100$ patients for validation. When validating computable phenotypes, simple random sampling (SRS) is often used. Under SRS, $100$ patients would be chosen for chart review with equal probability (Figure 3A). This strategy ignores all available information. Some phenotyping algorithms require an SRS validation study. However, for full-likelihood approaches like SMLEs, any information available for all patients can be used to design the validation study. Designing based on fully observed information ensures that the MAR assumption holds for the downstream analysis while allowing for many targeted strategies.
Figure 3: Diagram of partial validation studies with $N$ patients extracted from the electronic health records (EHR) and $n$ patients chosen for validation via chart review. The least informative strategy to select the validation study is A) simple random sampling. More informative targeted strategies include B) case-control sampling, C) balanced case-control sampling using discretized ALI, D) optimal stratified sampling using discretized ALI, and E) residual sampling based on the naive model. Within strata, the darker shaded areas indicate which patients would be selected.
When designing a validation study to promote statistical efficiency of the parameter estimates (i.e., to minimize the variance in the estimated logOR on ALI), we start at the end. Beginning with the planned logistic regression model of $Y$ given ($X$, $Z$), we work backward to determine which patients will be most informative. For this reason, the “best” designs vary across different estimators, with model-based ones, like SMLEs, offering the most design flexibility. Estimators that rely on the design to handle missingness, like inverse probability weighting (IPW), require nonzero sampling probability for all patients; model-based estimators do not.
We considered stratified and rank-based designs. Multiple validation waves were conducted to allow the protocol and study design to adapt as information accumulated, which is becoming common in this space [42, 34, 43]. Ultimately, simulations based on preliminary validation data guided our selection of the targeted design used in the application (see Section 2.6).
@@ 2.4.1 Stratified Designs
Given our binary outcome, the case-control (CC) design is a natural option, wherein $50$ patients each are chosen from subgroups with $Y=0$ and $Y=1$ (Figure 3B). The balanced case-control (BCC) design further stratifies on covariates, subsequently choosing equal (i.e., balanced) numbers from each stratum (Figure 3C). Continuous covariates must be discretized to form strata. For instance, we could use $X_{D}^{*}$, a discretized version of $X^{*}$, to define strata of patients with error-prone ALI above and below the average. Then, validate $25$ patients from each of the four ($Y$, $X_{D}^{*}$) strata. Importantly, we can still fit the model using the continuous values $X^{*}$.
The CC and BCC designs are intuitive and offer efficiency gains over SRS. Still, slightly more complicated designs are guaranteed to be theoretically “optimal;” they choose the number of patients from each stratum to achieve the lowest possible variance for the logOR on ALI [44, 45, 46, 47, 37, 48]. [43] proposed the optimal (OPT) design for model-based analysis of data with binary outcome and exposure misclassification (Figure 3D), which could apply here. However, we would achieve the lowest possible variance for $X_{D}^{*}$ rather than $X^{*}$. Increasing the number of unique values of $X_{D}^{*}$ (i.e., more strata) can help, but information will still be lost. Moreover, the OPT design depends on the true logORs, which are unknown; reliable placeholders for them are required to implement the design (often collected in a multi-wave framework).
@@ 2.4.2 Rank-Based Designs
Designs that rank patients rather than stratifying them offer improved granularity and sizable efficiency gains. For example, extreme tail sampling (ETS) designs rank patients by some value of interest and target those with the smallest or largest (i.e., most “extreme”) values. To identify patients most informative not just about ALI but its relationship to healthcare utilization, residual sampling (RS) is preferred [49]. To implement RS with measurement error, we begin by fitting the “naive” analysis model, i.e., using standard logistic regression procedures with $Y$, $X^{*}$ (in place of $X$), and $Z$ to obtain estimates for the logORs $\hat{\boldsymbol{\beta}}^{*}$ from the full sample of $1000$ patients. Then, the naive model residuals are computed for each patient $i$ as
@@
and used to rank patients. The largest residuals $r_{i}^{*}$ (in magnitude) correspond to patients whose observed outcome $Y_{i}$ had a low predicted probability (Figure 3E). The auditDesignR package contains R functions to sample from all of these designs [50].
@@ 2.5 Protocol: Enriched Validation Protocol to Reduce Errors and Missingness
Chart reviews were conducted over $6$ months by four clinical research assistants and technicians at WFUSM (hereafter, “auditors”), chosen for their familiarity with Epic (the institution’s electronic charting software) and demonstrated research experience.
@@ 2.5.1 Identifying Sources of Auxiliary Information
Before the chart reviews, our clinical collaborators developed the “audit roadmap” (Table 1). While it is unlikely that we would identify the original missing measurements, we harnessed clinical expertise to identify auxiliary information about missing discretized components. The roadmap suggested where auditors could look in Epic to locate supplemental information about missing values. It employs “anchors” to try and recover missing components using clinically-driven placeholders [15].
For example, if a patient were missing HbA1c, auditors would search their chart for a diagnosis of diabetes or impaired glycemic control. If either was present, the patient’s diagnosis would suggest that their missing HbA1c value would have exceeded the $6.5\%$ threshold. Thus, their validated ALI $X$ would include one point for HbA1c. If neither diagnosis was present, $X$ would continue to exclude this still-missing component.
Many components, like HbA1c, had multiple sources of auxiliary information, while others, like triglycerides, had just one. Only serum albumin, a blood test to measure liver and kidney function, did not have a supplemental diagnosis.
@@ 2.5.2 Validated Data Collection
In preparation for validation, the analytical team transformed the extracted EHR data from the original wide format (one row per patient encounter and columns for different variables) to a long format (one row per patient encounter per variable). This transformation streamlined the data structure for auditors: one value to review per row (Supplemental Figure S2).
Before sending the data to the auditors, the roadmap was incorporated, and an empty column was added for the “reviewed value” (i.e., the one in Epic). Depending on the scenario, auditors would enter different information for the reviewed value, and each data point was assigned one of five auditor findings (Figure 4).
@@ 2.6 Multi-Wave Strategy: Sampling Adaptively and Improving Tools
Choosing the “best” design often relies on information that is unknown at baseline, like EHR error rates or the magnitude of the association between ALI and healthcare utilization. Since we did not have any prior information about these unknowns, we started by sampling conservatively to collect preliminary information for a more extreme, targeted design later on. Multi-wave validation schemes like this one are appealing because they balance learning about our data with strategically designing for them [42, 34]. Even if prior information is available (e.g., from previous chart reviews), multi-wave validation allows us to relearn how data have changed before dedicating resources to validate them [43]. Further, we can seize opportunities to incorporate auditors’ feedback and improve the validation protocol.
@@
We validated $100$ patients’ data across three waves: the Pilot, Wave I, and Wave II. The small Pilot wave tested the validation protocol and established inter-auditor agreement on the first four patients. Then, Waves I and II put this protocol into action, as auditors validated the remaining patients. Waves I and II each selected $48$ more patients. (Designating equal-sized waves balances design flexibility and stability [42].) In Wave II, information from the Pilot + Wave I was leveraged to design a more targeted strategy. Patients could be chosen for validation at most once.
2.6.1 Adapting the Validation Study Design as Information Accumulates
@@
Between the Pilot and Wave I, the first $52$ patients were selected for chart review via the BCC design. Specifically, patients were split into four strata based on their outcome and error-prone ALI (discretized at the median of $0.33$). Then, equal numbers of patients were randomly chosen from the four strata. In the Pilot, one patient was chosen from each stratum; another $12$ were chosen per-stratum in Wave I.
@@
Before validating the remaining $48$ patients, we designed bespoke simulation studies based on the preliminary validation data. They explored how designs compared under different scenarios for true underlying data quality. The design offering the best efficiency (i.e., lowest variability) in estimating the logOR on ALI with the SMLE was desired.
@@
Under varied error rates, the targeted validation study designs offered empirically unbiased estimates and higher efficiency than SRS. The RS design provided the best efficiency under extra low to moderate error settings, and the preliminary validation data (Pilot + Wave I) indicated a perfect true positive rate (TPR) $=1.00$ and negligible false positive rate (FPR) $<0.01$. Under full/high data recovery rates ($90$-$100\%$ missing data recovered through validation) and different designs, the SMLE was empirically unbiased. While biased for moderate recovery or lower ($\leq 50\%$), the SMLE was still closer to the truth than the naive estimate was (i.e., based on unvalidated EHR data). Regardless of how much missing data was recovered, the RS design offered the highest efficiency. Details like the data-generating mechanisms can be found in the Supplemental Materials.
@@
We concluded that the RS design was the most informative way to select the remaining $48$ patients for validation in Wave II. To implement it, each unvalidated patient’s residual was calculated based on the naive healthcare utilization model estimates (Section 3.3) following (4), as
@@
Then, patients were ranked by $r_{i}^{*}$, and the $24$ patients with the smallest residuals and the $24$ patients with the largest residuals were selected for validation. Here, the RS design prioritized (i) patients with zero hospitalizations or ED visits but high predicted probabilities of having at least one and (ii) patients with at least one hospitalization or ED visit but high predicted probabilities of having zero. That is, patients whose observed $Y_{i}$ aligned well with the naive model’s prediction were not considered to be very informative and thus were deprioritized for validation.
@@ 2.6.2 Testing the Validation Protocol and Establishing Inter-Auditor Agreement
The validation protocol evolved slightly throughout the waves, as well. After completing the first four patients’ chart reviews (Pilot), auditors and analysts met to discuss any protocol modifications that would be helpful before proceeding. For example, the auditor finding “extracted value not found” was not one of the original options, and the team discussed how these data points should be recorded. Adding an optional “Notes” column, where auditors could include any additional information, also resulted from this discussion.
The Pilot also presented a valuable opportunity to evaluate inter-auditor agreement. We started with two auditors on our team. Each reviewed the four patients chosen for the Pilot, allowing us to calculate inter-auditor agreement in this doubly-validated sample. Later, two new auditors joined the team, each reviewing these four patients’ records as well. Ultimately, agreement in the four auditors’ findings was calculated using Fleiss’ Kappa [51], as detailed in the Supplemental Materials. We found high agreement between the auditors’ findings for labs ($\kappa=0.90$) and vitals ($\kappa=0.81$).
@@ 3.1 Cohort Description and Study Objectives
We considered adult patients ($18$-$65$ years old) with at least one primary care outpatient encounter at Atrium Health Wake Forest Baptist Hospital. The $1000$ sampled patients had a median age of $48$ years and a median income of \$52,320. They were $61\%$ female, $72\%$ White, and $6\%$ Hispanic. During the $2$-year study period, $318$ patients ($32\%$) had at least one hospitalization or ED visit. In unvalidated EHR data, error-prone ALI was right-skewed (Supplemental Figure S3). With a median of $0.33$ (interquartile range $=0.17,0.50$), patients experienced one out of every three non-missing components on their whole-person health, on average. We validated ALI data for $100$ patients.
First, we discuss the EHR data quality (Section 3.2). Second, we present the healthcare utilization model (Section 3.3). Within these sections, results are discussed chronologically, beginning with findings from the unvalidated EHR data and ending with those following the completion of all validation waves. We also discuss how these findings inform future implementations of the ALI in this healthcare system (Section 3.4).
@@ 3.2 Validation Study Findings: EHR Data Quality
As expected, the extracted EHR data had lots of missingness. Rare labs for inflammation components were missing for $>95\%$ of patients (Supplemental Figure S4A). Labs for metabolic components were missing for $21\%$ of patients. Common patterns of missing data are displayed in Supplemental Figure S5. Measurements for the cardiovascular components were non-missing for basically everyone. Fortunately, most patients ($78\%$) had at least five non-missing ALI components (Supplemental Figure S4B).
3.2.1 Pilot + Wave I Chart Reviews
@@
Figure 5: Auditors’ findings from the validation of A) $52$ patients’ data in the Pilot + Wave I, B) $48$ patients’ data in Wave II, and C) $100$ patients’ data in all waves combined. These findings refer to the original numeric measurements (before discretizing them into allostatic load index components), and there could be multiple per patient per variable. Auditor findings were mutually exclusive and specific to whether the measurements were originally non-missing (“extracted value correct,” “extracted value incorrect,” or “extracted value not found”) or missing (“no auxiliary information found” or “auxiliary information found”) in the extracted EHR data.
Figure 6: Comparison of the discretized allostatic load index components from the unvalidated EHR data to A) $52$ patients’ data in the Pilot + Wave I, B) $48$ patients’ data in Wave II, and C) $100$ patients’ data in all waves combined.
Fewer vitals than labs matched Epic ($59\%$ versus $81\%$), with $31\%$ not found. We discovered that the EHR data extraction algorithm expected every patient encounter to have vitals. For encounters where vitals were not taken (e.g., physical therapy), values were imputed from other encounters; imputed values could not be found in Epic by the auditors. Thus, we uncovered an area for potential improvement in the current EHR data extraction protocol.
Incorporating the audit roadmap into the validation protocol improved the completeness of the data. The number of missing values decreased for nearly all ALI components (Supplemental Figure S4A). Homocysteine saw the biggest improvement, with auxiliary information located for $23$ of $52$ validated patients ($44\%$). The number of non-missing ALI components per patient increased from $6.6$ to $7.5$, on average. Overall, $48$ of $177$ missing components ($27\%$) were recovered through validation (Figure 6A).

### v5

@@ 2.4 Design: Targeted Sampling to Select Informative Patients for Validation
Based on our analysis plan, we considered various sampling designs to select the most informative $100$ patients for validation. When validating computable phenotypes, simple random sampling (SRS) is often used. Under SRS, $100$ patients would be chosen for chart review with equal probability (Figure 3A). This strategy ignores all available information. Some phenotyping algorithms require an SRS validation study. However, for full-likelihood approaches like SMLEs, any information available for all patients can be used to design the validation study. Designing based on fully-observed information ensures that the MAR assumption holds for the downstream analysis while allowing for many targeted strategies.
Figure 3: Illustrative diagram of partial validation studies with $N=48$ patients extracted from the electronic health records (EHR) and $n=16$ patients chosen for validation via chart review. The least informative strategy to select the validation study is A) simple random sampling. More informative targeted strategies include B) case-control sampling, C) balanced case-control sampling using discretized ALI, D) optimal stratified sampling using discretized ALI, and E) residual sampling based on the naive healthcare utilization model. Within strata, the darker shaded areas indicate which patients would be selected.
When designing a validation study to promote statistical efficiency of a parameter estimate (e.g., to minimize the variance of the estimated logOR $\hat{\beta}_{1}$ on ALI), we start at the end. Beginning with the planned logistic regression model of $Y$ given ($X$, $Z$), we work backward to determine which patients will be most informative. For this reason, the “best” designs vary across different estimators, with model-based ones (like the SMLEs) offering the most flexibility. Estimators that rely on the design to handle missingness (like inverse probability weighting [IPW]) require a nonzero sampling probability for all patients; model-based estimators do not.
We considered stratified and rank-based designs. Multiple validation waves were conducted to allow the protocol and study design to adapt as information accumulated, which is becoming common practice in this space [42, 43, 35]. Ultimately, simulations based on preliminary validation data (Section 2.6) guided our selection of the targeted design used in the application.
@@ 2.4.1 Stratified Designs
Given our binary outcome, the case-control (CC) design is a natural option, wherein $50$ patients each are chosen with $Y=0$ and $Y=1$ (Figure 3B). The balanced case-control (BCC) design further stratifies on covariates and chooses equal (balanced) numbers from each stratum (Figure 3C). Continuous covariates must be discretized to form strata. For instance, we could let $X_{D}^{*}$ define strata of patients with error-prone ALI $X^{*}$ above and below the average and then validate $25$ patients from each of the four ($Y$, $X_{D}^{*}$) strata. Importantly, we can still fit the model using the continuous values $X^{*}$.
The CC and BCC designs are intuitive and offer efficiency gains over SRS. Still, slightly more complicated designs are guaranteed to be theoretically “optimal;” they choose the number of patients per stratum to achieve the lowest possible variance for $\hat{\beta}_{1}$ [44, 45, 46, 47, 48, 36]. Lotspeich et al. (2024) [42] proposed the optimal (OPT) design for model-based analyses with binary outcome and exposure misclassification (Figure 3D), which could apply here. However, we would achieve the lowest possible variance for $X_{D}^{*}$ rather than $X^{*}$. Increasing the number of unique values of $X_{D}^{*}$ (i.e., creating more strata) can help, but information will still be lost. Moreover, the OPT design depends on the true logORs $\boldsymbol{\beta}$}, which are unknown. Reliable placeholders for $\boldsymbol{\beta}$} (often collected in a multi-wave framework) are required to implement the design.
@@ 2.4.2 Rank-Based Designs
Designs ranking patients offer improved granularity and efficiency gains. For example, extreme tail sampling (ETS) designs order patients by some value of interest and target those with the smallest or largest (most “extreme”) values. To identify patients who are most informative to the healthcare utilization model, residual sampling (RS) is preferred [49]. We begin by fitting the “naive” analysis model, i.e., using standard logistic regression procedures with $Y$, $X^{*}$ (instead of $X$), and $Z$ to estimate the logORs $\hat{\boldsymbol{\beta}}^{*}$ from the full sample. Then, we compute residuals for all patients as
@@
and sample the $100$ patients with the most extreme residuals (Figure 3E). The largest residuals (in magnitude) correspond to patients whose observed $Y_{i}$ had a low predicted probability. The auditDesignR package contains R functions to sample from all of these designs [50].
@@ 2.5 Protocol: Enriched Validation Protocol to Reduce Errors and Missingness
Chart reviews were conducted over $6$ months by four clinical research technicians at WFUSM (hereafter, “auditors”), who were chosen for their familiarity with Epic (the institution’s electronic charting software) and research experience.
@@ 2.5.1 Identifying Sources of Auxiliary Information
Before the chart reviews, we developed the “audit roadmap” (Table 1), harnessing clinical expertise to identify auxiliary information about missing ALI components. The roadmap suggested where auditors could look in Epic to locate supplemental information about missing values. It employs “anchors” to try to recover missing components using clinically-driven placeholders [15].
For example, if a patient was missing HbA1c, auditors would search their chart for a diabetes or impaired glycemic control diagnosis. If either was present, the patient’s diagnosis would suggest that their missing HbA1c value would have exceeded the $6.5\%$ threshold. Thus, their validated $X$ would include one point for HbA1c. If neither diagnosis was present, $X$ would exclude this still-missing component.
Many components, like HbA1c, had multiple sources of auxiliary information, while others, like triglycerides, had just one. Only serum albumin, a blood test for liver and kidney function, did not have any.
@@ 2.5.2 Validated Data Collection
In preparation, the extracted EHR data were transformed from a wide (one row per patient encounter) to a long format (one row per patient encounter per variable). This transformation streamlined the data structure for auditors, with one value per row to review (Supplemental Figure S2).
Before sending the data to the auditors, we incorporated the roadmap and added an empty column for the “reviewed value” (i.e., the one in Epic). Depending on the scenario, auditors would enter different information for the reviewed value, and each data point was assigned one of five auditor findings (Figure 4).
@@ 2.6 Multi-Wave Strategy: Sampling Adaptively and Improving Tools
Choosing the “best” design often relies on information that is unknown at baseline, like the EHR error rates or the logORs for the healthcare utilization model. Since we did not have any prior information about these unknowns, we started by sampling conservatively to collect information for a targeted design later. This multi-wave validation scheme balances learning about our data with strategically designing for them. Even if prior information is available (e.g., from previous chart reviews), multi-wave validation allows us to learn how our data have changed before dedicating resources to validate them again [42]. Further, we can incorporate auditors’ feedback between waves to improve the validation protocol.
@@
We validated $100$ patients’ data across three waves: the Pilot, Wave I, and Wave II. The Pilot wave tested the validation protocol and established inter-auditor agreement on the first four patients. Then, Waves I and II put this protocol into action as auditors validated the remaining patients. Waves I and II each selected $48$ more patients. (Designating equal-sized waves balances design flexibility and stability [43].) In Wave II, information from the Pilot and Wave I was leveraged to design a more targeted strategy. Patients could be chosen for validation at most once.
2.6.1 Modifying the Validation Study Design as Information Accumulates
@@
Between the Pilot and Wave I, $52$ patients were selected for chart review via the BCC design. Specifically, patients were split into four strata based on their outcome and error-prone ALI (discretized at the median of $0.33$). Then, equal numbers of patients were randomly chosen from the four strata. In the Pilot, one patient was chosen from each stratum; in Wave I, another $12$ were.
@@
Before beginning Wave II, we designed simulation studies based on the preliminary validation data to explore how designs compared under different scenarios for true underlying data quality. The design offering the best efficiency (i.e., lowest variability) for the SMLE $\hat{\beta}_{1}$ of the logOR on ALI was desired. Details are in the Supplemental Materials.
@@
Under varied error rates, the targeted designs offered empirically unbiased estimates and higher efficiency than SRS. The RS design provided the best efficiency under extra-low to moderate error settings, and the preliminary validation data (Pilot and Wave I) indicated a true positive rate (TPR) =1.00=1.00 and a false positive rate (FPR) <0.01<0.01.
[…]
@@
We concluded that the RS design was the most informative way to select the remaining $48$ patients for validation in Wave II. Each unvalidated patient’s residual was calculated based on the naive healthcare utilization model estimates (Section 3.3) following (4) as
@@
Then, patients were ranked by $r_{i}^{*}$. The $24$ patients with the smallest residuals and the $24$ patients with the largest residuals were selected for validation. Here, the RS design prioritized (i) patients with zero hospitalizations or ED visits but high predicted probabilities of having at least one and (ii) patients with at least one hospitalization or ED visit but high predicted probabilities of having zero. That is, patients whose $Y_{i}$ aligned with the naive model’s prediction were not considered to be very informative and were deprioritized for validation.
@@ 2.6.2 Testing the Validation Protocol and Establishing Inter-Auditor Agreement
The validation protocol evolved slightly throughout the waves, as well. After completing the Pilot, auditors and analysts met to discuss any protocol modifications that would be helpful. For example, the auditor finding “extracted value not found” was not originally an option, and we discussed how these data points should be recorded. Adding an optional auditor “Notes” column resulted from this conversation, too.
We also evaluated inter-auditor agreement in the Pilot. We started with two auditors on our team. Each off them reviewed the four patients chosen for the Pilot, allowing us to calculate inter-auditor agreement in this doubly-validated sample. Later, two new auditors joined and also reviewed these four patients’ records. Ultimately, agreement in the four auditors’ findings was calculated using Fleiss’ Kappa [51], as detailed in the Supplemental Materials. We found high agreement for labs ($\kappa=0.90$) and vitals ($\kappa=0.81$).
@@ 3.1 Cohort Description and Study Objectives
We considered adult patients (18186565 years old) with at least one primary care outpatient encounter at Atrium Health Wake Forest Baptist Hospital. The $1000$ sampled patients had a median age of $48$ years and a median income of \$52,320. They were $61\%$ female, $72\%$ White, and $6\%$ Hispanic. During the $2$-year study period, $318$ patients ($32\%$) had at least one hospitalization or ED visit. In the unvalidated EHR data, error-prone ALI was right-skewed (Supplemental Figure S3). With a median of $0.33$ (interquartile range =0.17,0.50=0.17,0.50), patients experienced one out of every three non-missing components, on average. We validated ALI data for $100$ patients.
First, we discuss the EHR data quality. Second, we present the healthcare utilization model. Within these sections, results are introduced chronologically, beginning with findings from the unvalidated EHR data and ending with those based on all validation waves. We also explore how these findings inform future implementations of the ALI in this healthcare system.
@@ 3.2 Validation Study Findings: EHR Data Quality
As expected, the extracted EHR data had lots of missingness (Supplemental Figure S4). Rare inflammation labs were missing for $95\%$ of patients (Supplemental Figure S5A). Metabolic labs were missing for $21\%$ of patients. Cardiovascular components (from vitals) were non-missing for basically everyone. Fortunately, most patients ($78\%$) had at least five non-missing ALI components (Supplemental Figure S5B).
3.2.1 Pilot and Wave I Chart Reviews
@@
Figure 5: Auditors’ findings from the validation of A) $52$ patients’ data in the Pilot and Wave I, B) $48$ patients’ data in Wave II, and C) $100$ patients’ data in all waves combined. These findings refer to the original numeric measurements (before discretizing them into allostatic load index components), and there could be multiple per patient per variable. Auditor findings were mutually exclusive and specific to whether the measurements were originally non-missing (“extracted value correct,” “extracted value incorrect,” or “extracted value not found”) or missing (“no auxiliary information found” or “auxiliary information found”) in the extracted electronic health records data.
Figure 6: Comparison of the discretized allostatic load index components from the unvalidated electronic health records (EHR) data to A) $52$ patients’ data in the Pilot and Wave I, B) $48$ patients’ data in Wave II, and C) $100$ patients’ data in all waves combined.
Fewer vitals than labs matched Epic ($59\%$ versus $81\%$), with $31\%$ not found. We discovered that the EHR data extraction algorithm expected every patient encounter to have vitals. When vitals were not taken (e.g., physical therapy), values were imputed from other encounters; auditors could not find these values in Epic. Thus, we uncovered an improvement area in the EHR data extraction protocol.
Incorporating the roadmap into the validation protocol improved the completeness of the data. The number of missing values decreased for nearly all ALI components (Supplemental Figure S5A). Homocysteine saw the biggest improvement, with auxiliary information located for $23$ of $52$ validated patients ($44\%$). The median number of non-missing ALI components per patient increased from $7$ to $8$. Overall, $48$ of $177$ missing components ($27\%$) were recovered (Figure 6A).